OpenStudy (anonymous):

Solve the equation. Check your answer. square root x+7=5

OpenStudy (radar):

\[\sqrt{x+7}=5\] Square both sides. x+7 = 25 x=25-7=18

OpenStudy (radar):

Requires answer check. \[\sqrt{18+7}=5\]5=5 using only the positive root.